# Conformal Prediction with Corrupted Labels: Uncertain Imputation and Robust Re-weighting

### v2

@@ 4.2 Synthetic experiment: weights are hard to estimate
To demonstrate the advantages of UI compared to PCP, we examine a setup in which the corruption probabilities are hard to estimate, whereas $Z$ is a strong predictor of $Y$. We compare the performance of UI and PCP (applied with estimated weights) on the synthetic dataset from Section 4.1, with a different missing mechanism that is engineered to be challenging to estimate; see Appendix D.3 for details. Figure 4 shows that PCP fails to attain the nominal 90% coverage rate due to inaccuracies in the estimated weights. In contrast, UI relies on accurate estimates of $Y$ from $(X,Z)$ rather than on weights, therefore, it consistently achieves the desired coverage, as guaranteed by Theorem 4. This experiment demonstrates that UI can construct valid prediction sets, even in cases where PCP does not achieve the desired coverage rate.
Figure 4: Complex weights experiment. The coverage rate and average interval length obtained by Naive CP, PCP and UI. Performance metrics are evaluated over 30 random data splits.
@@ 4.3 Real datasets with artificial corruptions
In this section, we evaluate the proposed UI in a missing response setup using five benchmarks used in [5]: Facebook1,2 [22], Bio [23], House [24], Meps19 [2]. These datasets do not originally include privileged information, so we artificially define the PI as the feature in $X$ with the highest correlation to $Y$ and remove it from $X$, so that the PI is unavailable at test time. Since all true labels are available in the original datasets, we artificially remove 20% of them. The missing mechanism is intentionally designed to induce a distribution shift between missing and observed variables [5].
We employ the naive conformal prediction (Naive CP) that uses only the observed labels, PCP, which uses either the estimated corruption probabilities or the true ones when computing the weights $w(z)$, the proposed UI. Additionally, to demonstrate the importance of the error sampling scheme of UI, we apply CP with a naive imputation scheme, in which missing labels are replaced with the mean estimates of $Y\mid X,Z$ (Naive Imputation). The performances of all calibration schemes are presented in Figure 5. This figure indicates that the naive approaches produce too narrow intervals that do not achieve the desired coverage level. This is anticipated, as the validity guarantees of CP do not hold under distribution shifts or naive imputations. However, PCP, and the proposed UI consistently achieve the target 90% coverage level, as they appropriately account for the distributional shift. Notably, since PCP with estimated weights is valid while Naive CP is not, Figure 2 suggests that weight estimation errors fall within the theoretical validity region. Overall, this experiment reveals that UI constructs uncertainty intervals that are both statistically efficient and reliable.
Figure 5: Missing response experiment. The coverage rate and average interval length obtained by various methods; see text for details. Performance metrics are evaluated over 30 random data splits.
@@ 5 Discussion and impact statement
In this work, we analyzed the impact of inaccurate weights on WCP and PCP, and introduced UI, a novel calibration technique for reliable uncertainty quantification under corrupted labels. Our theoretical guarantees, supported by empirical experiments, demonstrate that UI constructs valid uncertainty estimates. While the validity conditions we derived for WCP and PCP are theoretically grounded, they require access to true weights, which are unavailable in practice. This calls for a promising future research direction in estimating these conditions from the available data. One limitation of our proposed UI is that it requires the features and responses to be independent of the corruption indicator given the privileged information, which resembles the strong ignorability assumption in causal inference [25, 26, 27]. Furthermore, our approach assumes that the label variability depends only on the PI and that the labels can be accurately estimated from the features and PI. In practice, our experiments indicate that our method attains valid coverage. Finally, we acknowledge potential social implications, akin to many developments in ML.
Acknowledgments and Disclosure of Funding
Y. R. and S. F. were supported by the ISRAEL SCIENCE FOUNDATION (grant No. 729/21). Y. R. and S. F. also thank the European Union (ERC, SafetyBounds, 101163414) for providing additional support. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Y.R. thanks the Career Advancement Fellowship, Technion.
@@ A.1.1 Notations
For ease of notations, we define $S_{n+1}=\infty$, while the score of the test sample is denoted by $\mathcal{S}(X_{n+1},Y_{n+1};\hat{f})$. Recall that the ground truth weights used by WCP are formulated by:
@@ A.1.2 Constant error
$\hat{p}_{i}:=\frac{\hat{w}_{i}}{\sum_{j=1}^{n}{\hat{w}_{j}}}=\frac{{w}_{i}+\delta}{\sum_{j=1}^{n}{{w}_{j}+\delta}}=\frac{{w}_{i}+\delta}{{W}_{n+1}+(n+1)\delta}.$ (26)
We begin with a general lemma that provides a deterministic connection between $k^{\texttt{CP}}$ and $k^{\texttt{WCP}}$, as defined in (22) and (21), respectively.
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta$
@@
We now turn to $-\frac{{W}_{n+1}}{n+1}<\delta<0$:
@@
Suppose that the calibration set is fixed, and equals to $\{(X_{i},Y_{i},Z_{i})=(x_{i},y_{i},z_{i})\}_{i=1}^{n}$ for some $x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},z_{i}\in\mathcal{Z}$ for all $i\in\{1,\ldots,n\}$. Further suppose that the test PI is fixed as well, i.e., $Z_{n+1}=z$ for some $z\in\mathcal{Z}$. Then, $\hat{C}^{\texttt{WCP}}(X_{n+1})$, as defined in (24), achieves a conservative coverage rate, i.e,:
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta$
@@
The probabilities in this proof are taken for drawing $(X_{n+1},Y_{n+1})\sim P_{X,Y}$ conditional on the fixed calibration set $\{(X_{i},Y_{i})=(x_{i},y_{i})\}_{i=1}^{n}$ and the fixed test PI $Z_{n+1}=z$. For ease of notation, we omit the conditioning on the calibration set in the formulas. Notice that in this case, $\hat{Q}^{\texttt{WCP}}$,$Q^{\texttt{WCP}}$, $Q^{\texttt{CP}}$ are deterministic since these are functions of the calibration set and the test PI $Z_{n+1}$. Observe that by the construction of the uncertainty sets:
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $-\frac{{W}_{n+1}}{n+1}<\delta$
@@
According to [5, Theorem 1], we get that when applied with the same weights, PCP achieves a higher coverage rate than WCP, that is:
@@
Lastly, under the exchangeability of the samples, [3, Corollary 1] states that:
@@ A.1.3 General bounded error
$\begin{split}&(a\text{ XOR }b)=((\text{not }a)\text{ and }b)\text{ or }(a\text{ and }(\text{not }b)),\\ &(a\text{ NXOR }b)=(\text{not }(a\text{ XOR }b)).\end{split}$ (51)
@@
$\begin{split}({W}_{k^{\texttt{WCP}}}+\Delta_{k^{\texttt{WCP}}}){W}_{n+1}&\leq({W}_{n+1}+\Delta_{n+1}){W}_{k^{\texttt{WCP}}}\iff\\ {\bar{\delta}}[{W}_{n+1}k^{\texttt{WCP}}-(n+1){W}_{k^{\texttt{WCP}}}]&\leq\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}\iff\\ 0&\leq\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}\iff\\ \tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}&\leq\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}\iff\\ \frac{\tilde{\Delta}_{k^{\texttt{WCP}}}}{\tilde{\Delta}_{n+1}}&\leq\frac{{W}_{k^{\texttt{WCP}}}}{{W}_{n+1}}\\ \end{split}$ (59)
@@
Suppose that the calibration set is fixed, and equals to $\{(X_{i},Y_{i},Z_{i})=(x_{i},y_{i},z_{i})\}_{i=1}^{n}$ for some $x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},z_{i}\in\mathcal{Z}$ for all $i\in\{1,\ldots,n\}$. Further suppose that the test PI and the errors $\{\delta_{i}\}_{i=1}^{n+1}$ are also fixed, i.e., $Z_{n+1}=z$ for some $z\in\mathcal{Z}$. Then, $\hat{C}^{\texttt{WCP}}(X_{n+1})$, as defined in (24), achieves a conservative coverage rate, i.e,
@@
    $k^{\texttt{CP}}>k^{\texttt{WCP}}$ and (requirement 1 NXOR requirement 2)
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and (requirement 1 XOR requirement 2)
@@
    $\mathbb{P}\left.(k^{\texttt{CP}}>k^{\texttt{WCP}},\text{(requirement 1 NXOR requirement 2)}\right.)\geq 1-\varepsilon$,
@@
    $\mathbb{P}\left.(k^{\texttt{CP}}<k^{\texttt{WCP}},\text{(requirement 1 XOR requirement 2)}\right.)\geq 1-\varepsilon$,
@@ A.2 Uncertain imputation
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)+v)\geq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)-v)$, and
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)-v)\geq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)+v)$.
@@
[…]
We denote by $R$ the sum of $R^{\text{test}}$ and $-D^{\prime}$:
@@
[…]
@@
[…]
@@
[…]
@@
$\bar{Y}^{\text{test}}:=\begin{cases}Y^{\text{test}},&M^{\text{test}}=0,\\ \hat{g}(X^{\text{test}},Z^{\text{test}})+e,&M^{\text{test}}=1,\\ \end{cases}$ (85)
@@
$\bar{Y}^{\text{test}}:=\begin{cases}Y^{\text{test}},&M^{\text{test}}=0,\\ \hat{g}(X^{\text{test}},Z^{\text{test}})+e,&M^{\text{test}}=1.\\ \end{cases}$ (94)
Above, $e$ is a random variable drawn from the distribution of $Y-\hat{g}(X,Z)\mid Z=Z^{\text{test}},M=0$. Since $(X,Y)\perp\!\!\!\perp M\mid Z$, this distribution is equivalent to the distribution of $Y-\hat{g}(X,Z)\mid Z=Z^{\text{test}}$. Since $C^{\texttt{UI}}$ is constructed by CP using the imputed labels, and due to the exchangeability of the data, $C^{\texttt{UI}}$ covers $\bar{Y}^{\text{test}}$ at the desired coverage rate [1, 28]:

### v3

@@ 4.2 Synthetic experiment: weights are hard to estimate
To demonstrate the advantages of UI compared to PCP, we adopt the synthetic setup from Feldman & Romano (2024), where $Z$ is a strong predictor of $Y$. The only difference is that we engineer the missingness mechanism to be challenging to estimate; see Appendix D.3 for details. Figure 4 shows that PCP fails to attain the nominal 90% coverage rate due to inaccuracies in the estimated weights. In contrast, UI relies on accurate estimates of $Y$ from $(X,Z)$ rather than on weights; therefore, it consistently achieves the desired coverage, as guaranteed by Theorem 4. This experiment demonstrates that UI can achieve the desired coverage rate, even in cases where PCP does not.
Figure 4: Complex weights experiment. The performance of Naive CP, PCP and UI.
@@ 4.3 Real datasets with artificial corruptions
In this section, we evaluate the proposed UI in a missing response setup using five benchmarks used in Feldman & Romano (2024): Facebook1,2 (facebook, ), Bio (bio, ), House (house, ), Meps19 (meps_, 19). We follow Feldman & Romano (2024) and artificially define the PI as the feature in $X$ with the highest correlation to $Y$ and remove it from $X$, so that the PI is unavailable at test time. Since all true labels are available in the original datasets, we artificially remove 20% of them similarly to Feldman & Romano (2024) to induce a distribution shift between missing and observed variables.
We employ the naive conformal prediction (Naive CP) that uses only the observed labels, PCP, with either estimated corruption probabilities or true ones used for computing the weights $w(z)$, and the proposed UI. Additionally, to demonstrate the importance of the error sampling scheme of UI, we apply CP with a naive imputation scheme that replaces missing labels with the mean estimates of $Y\mid X,Z$ (Naive Imputation). The performances of all calibration schemes are presented in Figure 5. This figure indicates that the naive approaches produce too narrow intervals that do not achieve the desired coverage level. This is anticipated, as the validity guarantees of CP do not hold under distribution shifts or naive imputations. However, PCP, and the proposed UI consistently achieve the target 90% coverage level, as they appropriately account for the distributional shift. This is also indicated by Theorem 4. Notably, since PCP with estimated weights is valid while Naive CP is not, Figure 2 suggests that weight estimation errors fall within the theoretical validity region. Overall, this experiment reveals that UI constructs uncertainty intervals that are both statistically efficient and reliable. The performance of TriplyRobust in this experiment is provided in Appendix F.6.
Figure 5: Missing response experiment. The performance of various methods; see text for details.
@@ 5 Discussion and impact statement
In this work, we analyzed the impact of inaccurate weights on WCP and PCP, and introduced UI, a novel calibration technique for reliable uncertainty quantification under corrupted labels. Our theoretical guarantees, supported by empirical experiments, demonstrate that UI constructs valid uncertainty estimates. While the validity conditions we derived for WCP and PCP are theoretically grounded, they require access to true weights, which are unavailable in practice. This calls for a promising future research direction in estimating these conditions from the available data. One limitation of our proposed UI is that it requires the features and responses to be independent of the corruption indicator given the privileged information, which resembles the strong ignorability assumption in causal inference (Rubin, 1978; Rosenbaum & Rubin, 1983; Imbens & Rubin, 2015). Furthermore, our approach assumes that the label variability depends only on the PI and that the labels can be accurately estimated from the features and PI. In practice, our experiments indicate that our method attains valid coverage. Future work could explore extending our theoretical guarantees to multiple-annotator settings and incorporating privileged information into ambiguity-aware calibration methods Javanmardi et al. (2024); Caprio et al. (2025). Finally, we acknowledge potential social implications, akin to many developments in ML.
Acknowledgments
Funded by the European Union (ERC, SafetyBounds, 101163414). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. Y.R. thanks the Career Advancement Fellowship, Technion.
@@ A.1.1 Notations
[…]
@@ A.1.2 Constant error
$\hat{p}_{i}:=\frac{\hat{w}_{i}}{\sum_{j=1}^{n+1}{\hat{w}_{j}}}=\frac{{w}_{i}+\delta}{\sum_{j=1}^{n+1}{({w}_{j}+\delta})}=\frac{{w}_{i}+\delta}{{W}_{n+1}+(n+1)\delta}.$ (22)
We begin with a general lemma that provides a deterministic connection between $k^{\texttt{CP}}$ and $k^{\texttt{WCP}}$, as defined in equation 18 and equation 17, respectively.
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta$
@@
We now turn to $\frac{{W}_{n+1}}{n+1}<\delta<0$:
@@
Suppose that the calibration set is fixed, and equals to $\{(X_{i},Y_{i},Z_{i})=(x_{i},y_{i},z_{i})\}_{i=1}^{n}$ for some $x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},z_{i}\in\mathcal{Z}$ for all $i\in\{1,\ldots,n\}$. Further suppose that the test PI is fixed as well, i.e., $Z_{n+1}=z$ for some $z\in\mathcal{Z}$. Then, $\hat{C}^{\texttt{WCP}}(X_{n+1})$, as defined in equation 20, achieves a conservative coverage rate, i.e,:
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta$
@@
The probabilities in this proof are taken for drawing $(X_{n+1},Y_{n+1})\sim P_{X,Y}$ conditional on the fixed calibration set $\{(X_{i},Y_{i})=(x_{i},y_{i})\}_{i=1}^{n}$ and the fixed test PI $Z_{n+1}=z$. For ease of notation, we omit the conditioning on the calibration set in the formulas. Notice that in this case, $\hat{Q}^{\texttt{WCP}}$$Q^{\texttt{WCP}}$, $Q^{\texttt{CP}}$ are deterministic since these are functions of the calibration set and the test PI $Z_{n+1}$. Observe that by the construction of the uncertainty sets:
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta\leq 0$
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and $\frac{{W}_{n+1}}{n+1}<\delta$
@@
According to Feldman & Romano (2024, Theorem 1), we get that when applied with the same weights, PCP achieves a higher coverage rate than WCP, that is:
@@
Lastly, under the exchangeability of the samples, Tibshirani et al. (2019, Corollary 1) states that:
@@ A.1.3 General bounded error
[…]
@@
[…]
@@
Suppose that the calibration set is fixed, and equals to $\{(X_{i},Y_{i},Z_{i})=(x_{i},y_{i},z_{i})\}_{i=1}^{n}$ for some $x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},z_{i}\in\mathcal{Z}$ for all $i\in\{1,\ldots,n\}$. Further suppose that the test PI and the errors $\{\delta_{i}\}_{i=1}^{n+1}$ are also fixed, i.e., $Z_{n+1}=z$ for some $z\in\mathcal{Z}$. Then, $\hat{C}^{\texttt{WCP}}(X_{n+1})$, as defined in equation 20, achieves a conservative coverage rate, i.e,
@@
    $k^{\texttt{CP}}<k^{\texttt{WCP}}$ and (requirement 1 NXOR requirement 2)
@@
    $k^{\texttt{CP}}>k^{\texttt{WCP}}$ and (requirement 1 XOR requirement 2)
@@
    $\mathbb{P}\left.(k^{\texttt{CP}}<k^{\texttt{WCP}},\text{(requirement 1 NXOR requirement 2)}\right.)\geq 1-\varepsilon$,
@@
    $\mathbb{P}\left.(k^{\texttt{CP}}>k^{\texttt{WCP}},\text{(requirement 1 XOR requirement 2)}\right.)\geq 1-\varepsilon$,
@@ A.2 Uncertain imputation
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)+v)\leq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)-v)$, and
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)-v)\leq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)+v)$.
@@
is equal in distribution to $Y^{\text{test}}$, that is: $Y^{\prime}\overset{d}{=}Y^{\text{test}}$. Furthermore, since $\varepsilon^{\text{test}}\perp\!\!\!\perp X^{\text{test}}$ we also get that:
$Y^{\prime}\mid X^{\text{test}}\overset{d}{=}Y^{\text{test}}\mid X^{\text{test}}.$ (66)
We denote by $R$ the sum of $R^{\text{test}}$ and $D^{\prime}$:
@@
Observe that since $R^{\text{test}}\overset{d}{=}D^{\prime}$ we get that $R$ is a symmetric random variable:
@@
$\begin{split}\mathbb{P}(\bar{Y}^{\text{test}}\leq B)&=\mathbb{P}(Y^{\prime}+R\leq B)\\ &=\mathbb{P}({g}^{*}(X^{\text{test}},Z^{\text{test}})+\varepsilon+R\leq B)\\ &=\mathbb{P}({g}^{*}(X^{\text{test}},Z^{\text{test}})+\varepsilon^{\text{test}}+R\leq B)\\ &=\mathbb{P}(Y^{\text{test}}+R\leq B)\\ &=\mathbb{P}(Y^{\text{test}}+R\leq B\mid R\geq 0)\mathbb{P}(R\geq 0)+\mathbb{P}(Y^{\text{test}}+R\leq B\mid R<0)\mathbb{P}(R<0)\\ &=\frac{1}{2}\left[\mathbb{P}(Y^{\text{test}}+R\leq B\mid R\geq 0)+\mathbb{P}(Y^{\text{test}}+R\leq B\mid R<0)\right]\\ &=\frac{1}{2}\left[\mathbb{P}(Y^{\text{test}}+R\leq B\mid R\geq 0)+\mathbb{P}(Y^{\text{test}}-R\leq B\mid R\geq 0)\right]\\ &=\frac{1}{2}\left[\mathbb{P}(Y^{\text{test}}\leq B-R\mid R\geq 0)+\mathbb{P}(Y^{\text{test}}\leq B+R\mid R\geq 0)\right]\\ &=\frac{1}{2}\int_{r\geq 0}\left[\mathbb{P}(Y^{\text{test}}\leq B-R\mid R=r)+\mathbb{P}(Y^{\text{test}}\leq B+R\mid R=r)\right]f_{R\mid R\geq 0}(r)dr.\end{split}$ (71)
@@
$\begin{split}\mathbb{P}(\bar{Y}^{\text{test}}\leq B)&=\frac{1}{2}\int_{r\geq 0}\left[\mathbb{P}(Y^{\text{test}}\leq B-R\mid R=r)\right.\\ &+\left.\mathbb{P}(Y^{\text{test}}\leq B+R\mid R=r)\right]f_{R\mid R\geq 0}(r)dr\\ &\leq\frac{1}{2}\int_{r\geq 0}\left[2\mathbb{P}(Y^{\text{test}}\leq B\mid R=r)\right]f_{R\mid R\geq 0}(r)dr\\ &=\mathbb{P}(Y^{\text{test}}\leq B\mid R\geq 0)\\ &=\mathbb{P}(Y^{\text{test}}\leq B).\end{split}$ (77)
@@
[…]
@@
[…]
Above, $e$ is a random variable drawn from the distribution of $Y-\hat{g}(X,Z)\mid Z=Z^{\text{test}},M=0$. Since $(X,Y)\perp\!\!\!\perp M\mid Z$, this distribution is equivalent to the distribution of $Y-\hat{g}(X,Z)\mid Z=Z^{\text{test}}$. Since $C^{\texttt{UI}}$ is constructed by CP using the imputed labels, and due to the exchangeability of the data, $C^{\texttt{UI}}$ covers $\bar{Y}^{\text{test}}$ at the desired coverage rate Vovk et al. (2005); Angelopoulos & Bates (2023):